A jet plane leaves Los Angeles for New York traveling for 5 hours with a wind pushing it. On the return trip the plane took 8 hours and there is still wind. What would be the speed of the plane if there wasn't any wind? How fast is the wind?
Can you please include the steps

Answer by richwmiller(17219)

Not enough info given
what is the distance from la to nyc?
assuming 2464 miles
5(p+w)=8(p-w)=2464
5p+5w=8p-8w=2464
p = 2002/5, w = 462/5